# A shared-revenue Bertrand game

### v1

@@ F.4. Upper bounding the retailer’s payoff
Analogous to the previous subsection, we know that the upper bound for the retailer’s payoff will be with the strategy profile $\underline{\rho}\equiv\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\}$. The retailer’s maximum payoff either occurs when the independent seller fulfills demand at $p_{r^{*}}$ or $p_{s_{\text{ind}}}$, or when the retailer fulfills demand themselves at $p_{r^{*}}$. Recall that $p_{r^{*}}\leq p_{s_{\text{ind}}}\iff\alpha\leq\alpha_{r^{*}}$ per (46). The retailer’s payoff function when the independent seller fulfills at $p_{r^{*}}$ is simple –
$\pi_{r,s}(p_{r^{*}},\alpha)=\alpha p_{r^{*}}(1-p_{r^{*}})=\frac{\alpha}{4}(1-c_{r}^{2})$
This is clearly increasing in $\alpha$, so the maximum payoff that the retailer could achieve when the independent seller fulfills at $p_{r^{*}}$ is by setting $\alpha=\alpha_{r^{*}}$. At this $\alpha$, we have
(84) $\pi_{r,s}(p_{r^{*}},\alpha_{r^{*}})=\frac{1-2c_{s}+c_{r}}{4(1+c_{r})}\left(\frac{1-c_{r}}{2}\right)^{2}=\frac{1+c_{r}-2c_{s}}{4(1+c_{r})}\left(1-c_{r}^{2}\right)=\left(\frac{1+c_{r}-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$
Notice that this yields a higher payoff than the retailer fulfilling demand on their own, because
$\left(\frac{1+c_{r}-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$ $>\left(\frac{1-c_{r}}{2}\right)^{2}=\pi_{r,r}(p_{r^{*}})$
$1+c_{r}-2c_{s}$ $>1-c_{r}$
$c_{r}$ $>c_{s}$
Finally, it remains to see if the retailer could achieve a higher payoff than (84) for $\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]$ at price $p_{s_{\text{ind}}}$. The payoff function is slightly less simple –
$\pi_{r,s}(p_{s_{\text{ind}}},\alpha)=\alpha\left(\frac{c_{s}}{1-\alpha}\right)\left(1-\frac{c_{s}}{1-\alpha}\right)$
it’s not obvious if this is increasing or decreasing in $\alpha$, so we will take the derivative:
(85) $\frac{\partial\pi_{r,s}(p_{s_{\text{ind}}},\alpha)}{\partial\alpha}=\frac{\partial}{\partial\alpha}\left(\frac{\alpha c_{s}(1-\alpha-c_{s})}{(1-\alpha)^{2}}\right)=\frac{c_{s}[1-c_{s}-\alpha(1+c_{s})]}{(1-\alpha)^{3}}$
Clearly, the unique maximum occurs when $\alpha=\frac{1-c_{s}}{1+c_{s}}$, and $\pi_{r,s}(p_{s_{\text{ind}}},\alpha)$ is decreasing in $\alpha$ for $\alpha>\frac{1-c_{s}}{1+c_{s}}$. However, we have that
(86) $\frac{1-c_{s}}{1+c_{s}}$ $<1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
$\frac{2c_{s}}{1+c_{r}}$ $<\frac{2c_{s}}{1+c_{s}}$
$c_{s}$ $<c_{r}$
Thus, the maximum equilibrium payoff the retailer can attain when the retailer sells at $p_{s_{\text{ind}}}$ is at the minimum possible $\alpha$, which is in fact $\alpha=\alpha_{r^{*}}$. Conventiently, this implies that the maximizer of the retailer’s payoff function with the strategy $\underline{\rho}$ is precisely $\underline{\alpha}\equiv\alpha_{r^{*}}$. Because at $\underline{\alpha}$, we have $p_{s_{\text{ind}}}=p_{r^{*}}$, this is sufficient to conclude that
$\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)\leq\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\underline{\rho})=\pi_{r}^{\text{(eq)}}(\underline{\alpha},c_{r},c_{s},\underline{\rho})=\left(\frac{1+c_{r}-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$
@@
Just as the previous case, this is unfortunately not sufficient to conclude that $\alpha_{*}\leq\underline{\alpha}$ in general; the best we can do is conclude that $\alpha_{*}\leq\alpha^{\dagger}_{s}$. For the independent seller, it proves the trivial bound that $\pi_{s}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)\geq 0$, where the bound is tight because when $\alpha=\underline{\alpha}$ and $\rho=\underline{\rho}$, the independent seller is fulfilling demand at $p_{s_{\text{ind}}}$ and achieves a payoff of 00.
Appendix G Equilibria of leaving subgame
The leaving subgame is simply a standard Bertrand game with asymmetric costs $c_{s}+\delta<c_{r}$. Though this has been derived in the literature for general demand curves such as in (Blume, 2003; Kartik, 2011), for completeness we include here a brief derivation of the equilibria contextualized to our setting. First, note that $\beta\in\{0,1\}$, which follows as a corollary of Proposition 1 when $\alpha=0$. Thus, any shared-price equilibrium must have exactly one player fulfill demand. However, we have that
Lemma 0.
In all Nash equilibria of the leaving subgame, the independent seller fulfills demand.
Proof.
Suppose for contradiction there exists such an equilibrium $(p_{r},p_{s}\geq p_{r})$. Clearly, we must have $p_{r}\geq c_{r}$, otherwise at a price $p_{r}^{\prime}<c_{r}$ the retailer would prefer to increase their price to $p_{s}+\varepsilon$ and achieve a payoff of 00 rather than achieving negative payoff for each unit sold at $p_{r}^{\prime}$. However, this implies that $p_{s}\geq c_{r}>c_{s}+\delta$, so setting any price $p_{s}^{\prime}\in(c_{s}+\delta,c_{r})$ yields the independent seller a payoff of
$0=\pi_{s}^{(\ell)}(p_{r},p_{s})<\pi_{s}^{(\ell)}(p_{r},p_{s}^{\prime})=(p_{s}^{\prime}-(c_{s}+\delta))(1-p_{s}^{\prime})$
With this, consider an equilibrium $(p_{r}\geq p_{s},p_{s})$ in which the independent seller fulfills demand. In order for this to be an equilibrium, we must have that
$p_{s}=\arg\max_{p_{s}^{\prime}\in[0,p_{r}]}\pi_{s}^{(\ell)}(p_{r},p_{s}^{\prime})$
However, because $\pi_{s}^{(\ell)}$ a concave quadratic maximized at $p_{s^{*}}^{(\ell)}\equiv\frac{1+(c_{s}+\delta)}{2}$, it follows that the equilibrium $p_{s}$ is given by
$p_{s}=\begin{cases}p_{r}&p_{r}\leq p_{s^{*}}^{(\ell)}\\ p_{s^{*}}^{(\ell)}&p_{s^{*}}^{(\ell)}\leq p_{r}\end{cases}$
Consider the first case where $p_{r}\leq p_{s^{*}}^{(\ell)}$. If $p_{r}>p_{s}$, because $\pi_{s}^{(\ell)}$ is increasing for $p_{s}\in(0,p_{s^{*}})\supset(0,p_{r}]$, the independent seller achieves a higher payoff by setting price $p_{s}+\varepsilon$ for some $\varepsilon\in(0,p_{r}-p_{s})$. We therefore must have a shared price $p_{r}=p=p_{s}$ in this case. If $p>c_{r}$, the retailer achieves a higher price by setting $p^{\prime}\in(c_{r},p)$ and fulfilling some demand, yielding them a nonzero payoff. Furthermore, if $p>p_{s^{*}}^{(\ell)}$, the independent seller would prefer to fulfill at their optimal price $p_{s^{*}}^{(\ell)}$ instead. Thus, we have shared-price Nash equilibria at $p\in\left[c_{s}+\delta,\min\left\{p_{s^{*}}^{(\ell)},c_{r}\right\}\right]$.
Now consider the second case where $p_{s^{*}}^{(\ell)}\leq p_{r}$. If $c_{r}\leq p_{s^{*}}^{(\ell)}$, our analysis reduces to that of the previous case. If, on the other hand, $c_{r}>p_{s^{*}}^{(\ell)}$, then we still have similar shared-price equilibria $(p\in[c_{s}+\delta,p_{s^{*}}^{(\ell)}],s)$. However, we also gain unequal-price Nash equilibria $(p_{r}>p_{s},p_{s}=p_{s^{*}}^{(\ell)})$ – the independent seller would no longer benefit from deviating to a price $p_{s}+\varepsilon$ and the retailer still would not prefer to undercut the independent seller because $p_{s^{*}}^{(\ell)}<c_{r}$.
To summarize, we have found the following Nash equilibria:
$\begin{cases}p\in\left[c_{s}+\delta,c_{r}\right]&c_{r}\leq p_{s^{*}}^{(\ell)}\\ \left(p\in\left[c_{s}+\delta,p_{s^{*}}^{(\ell)}\right]\right)\cup\left(p_{r}>p_{s},p_{s}=p_{s^{*}}^{(\ell)}\right)&p_{s^{*}}^{(\ell)}\leq c_{r}\end{cases}$
Now that we have found the Nash equilibria, we apply the admissibility refinement that we have been using throughout. Very similar analysis to Appendix E.1 shows that the admissible strategy sets for the retailer and seller are $p_{r}\in[c_{r},p_{r^{*}}]$ and $p_{s}\in\left[c_{s}+\delta,p_{s^{*}}^{(\ell)}\right]$. While the independent seller is not playing any dominated strategies in equilibrium, the retailer is, so this is a nontrivial refinement. The admissible Nash equilibria are:
$(p_{r},p_{s})=\begin{cases}c_{r}&c_{r}\leq p_{s^{*}}^{(\ell)}\\ \left(p_{r}\in[c_{r},p_{r^{*}}],p_{s}=p_{s^{*}}^{(\ell)}\right)&p_{s^{*}}^{(\ell)}\leq c_{r}\end{cases}$
From here, we find that the admissible Nash equilibrium outcomes are:
$(p,x)=\begin{cases}(c_{r},s)&c_{r}\leq p_{s^{*}}^{(\ell)}\\ (p_{s^{*}}^{(\ell)},s)&p_{s^{*}}^{(\ell)}\leq c_{r}\end{cases}$
or simply $(p,x)=\left(\min\left\{c_{r},p_{s^{*}}^{(\ell)}\right\},s\right)$.

### v2

@@ F.4. Upper bounding the retailer’s payoff
Analogous to the previous subsection, we know that the upper bound for the retailer’s payoff will be with the strategy profile $\underline{\rho}\equiv\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\}$. The retailer’s maximum payoff either occurs when the independent seller fulfills demand at $p_{r^{*}}$ or $p_{s_{\text{ind}}}$, or when the retailer fulfills demand themselves at $p_{r^{*}}$.
Recall that $p_{s_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{r^{*}}$. Notice that $\pi_{r,s}(p_{r^{*}},\alpha)=\alpha p_{r^{*}}q(p_{r^{*}})$ is increasing in $\alpha$ because $p_{r^{*}}$ is independent of $\alpha$. Thus, it immediately follows that
(62) $\max_{\alpha\in\left[\max\left\{\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right\},\alpha_{r^{*}}\right]}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\underline{\rho})=\max_{\alpha\in\left[\max\left\{\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right\},\alpha_{r^{*}}\right]}\pi_{r,s}(p_{r^{*}},\alpha)=\pi_{r,s}(p_{r^{*}},\alpha_{r^{*}})$
At $\alpha_{r^{*}}=1-\frac{c_{s}}{p_{r^{*}}}$ per (36) we have
$\pi_{r,s}(p_{r^{*}},\alpha_{r^{*}})=\left(1-\frac{c_{s}}{p_{r^{*}}}\right)p_{r^{*}}q(p_{r^{*}})=(p_{r^{*}}-c_{s})q(p_{r^{*}})>(p_{r^{*}}-c_{r})q(p_{r^{*}})=\pi_{r,r}(p_{r^{*}})$
which implies the retailer always prefers to set $\alpha_{r^{*}}$ rather than any $\alpha\geq\alpha^{\dagger}_{s}$ where they fulfill demand themselves.
Now let’s turn our attention to the other half of the interval, $\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]$, on which the equilibrium is that the independent seller fulfills demand at $p_{s_{\text{ind}}}$. Indeed, on this interval we claim that $\pi_{r,s}(p_{s_{\text{ind}}},\alpha)$ is decreasing in $\alpha$. Indeed, notice $p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}$ is increasing in $\alpha$, and at $\alpha_{r^{*}}$, $p_{s_{\text{ind}}}=p_{r^{*}}>\tilde{p}$ per Lemma 2. Thus, for all $\alpha\geq\alpha_{r^{*}}$, we have $p_{s_{\text{ind}}}\geq p_{r^{*}}>\tilde{p}$, and since $\pi_{r,s}$ is unimodal per Corollary 3, this is sufficient to conclude the proof of the claim. In light of this, we have
(63) $\max_{\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\underline{\rho})=\max_{\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]}\pi_{r,s}(p_{s_{\text{ind}}},\alpha)=\pi_{r,s}(p_{s_{\text{ind}}},\alpha_{r^{*}})$
Combining this with (62) is sufficient to conclude that
$\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\underline{\rho})=\pi_{r}^{\text{(eq)}}(\alpha_{r^{*}},c_{r},c_{s},\underline{\rho})=\pi_{r,s}(p_{r^{*}},\alpha_{r^{*}})=\pi_{r,s}(p_{s_{\text{ind}}},\alpha_{r^{*}})$
@@
Just as the previous case, this is unfortunately not sufficient to conclude that $\alpha_{*}\leq\alpha_{r^{*}}$ in general; the best we can do is conclude that $\alpha_{*}\leq\alpha^{\dagger}_{s}$. For the independent seller, it results only in the trivial bound that $\pi_{s}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)\geq 0$, where the bound is tight because when $\alpha=\alpha_{r^{*}}$ and $\rho=\underline{\rho}$, the independent seller is fulfilling demand at $p_{s_{\text{ind}}}$ and achieves a payoff of 0.